# Self-organisation - the underlying principle and a general model

### v1

@@ VII. Concluding discussion
The principle underlying emergence of SO in out-of-equilibrium driven disordered systems has been formulated. It states that SO is the result of survivability of a minute subset of system configurations, or microstates, that stand out in their exceptional stability against the noise introduced by the driving. This principle is clearly supported by observations of SO in 2D granular systems, which evolve toward stable configurations of coordinated cell structures and stresses. Based on this principle, a statistical mechanical formalism has been constructed for SO in those 2D systems and its use outlined. Similar stability measures have been identified in 3D granular systems, which makes possible a straightforward extension to a similar model and its outline has been discussed in detail. The principle goes beyond granular systems - driven disordered systems always experience noisy dynamics and, when their steady states self-organise, they exclude the overwhelming majority of theoretically possible configurations which do not survive the noise. It also goes beyond the previously proposed principle of minimisation of dissipation energy Cyetal21 ; Keetal23 , which is only a special case that applies to systems where energy is the main driver of the SO. The generality of this principle makes possible an extension of the statistical mechanics approach to describe emergence of SO in wider range of systems. This extension has been formulated here explicitly and it involves identification of quasi-particles; their DOFs; stability measures, which that depend on the DOFs; and a survivability function, whose value decreases with the stability measures. The latter is used in a Boltzmann-like exponential, which determines the occurrence probability of system configurations, and a partition function is constructed. From the partition function, all expectation values of the self-organised state can be derived. This complements the idea in Keetal23 of the system exploring the configuration space in that it parks on configurations that minimise locally the survivability function. To illustrate the general formalism, it is applied to construct a statistical mechanics model for a very simple version of the laning problem.
@@
There are parallels between this formulation and its analogue in thermal systems, with system-wide configurations being the microstates and the stability measures paralleling the particles’ energies and potentials. The thermal survivability function is then the free energy, whose lowering increases a microstate’s occurrence probability and thence its survivability. Like the free energy, the survivability function depends on the quasi-particles’ DOFs. With this understanding on board, equilibrium states can also be viewed as self-organised. An example is solidification under temperature lowering, which is analogous to increasing $J_{i}$ in (4). the overwhelming number of microstates with very energetic particles are strongly unstable and only the minute crystalline microstates can be observed. There are, however, differences. The energy in thermal systems is a constant of the motion, from which dynamic equations for the DOFs can be derived. The survivability function only describes stability in out-of-equilibrium systems and, as far as this author knows, cannot play such a role. Additionally, thermal systems incur indigenous thermal fluctuations while in a-thermal systems it is the driving that gives rise to noise. Nevertheless, the similarities to traditional statistical mechanics are useful as guidelines for constructing general models of SO.
@@
One can argue that the above principle also applies to the SO that is evident in the evolution of biological species. Those systems are driven by many constraints: resources, predators, environmental, etc., which they must adapt to and survive. The dynamics are then extremely noisy because the constraints magnitudes and even number can fluctuate wildly. The principle of the survival of the fittest is exactly the same, with only individual elements of the systems that can survive these fluctuations, be it cells or individuals organisms, are realised. While this may mean that, in principle, it could be possible to formulate a statistical mechanical model for species evolution, it is a daunting task due to the enormous number of potential DOFs and stability measures.

### v2

@@ VII. Concluding discussion
The principle underlying emergence of SO in out-of-equilibrium driven disordered systems has been formulated. It states that SO is the result of survivability of a minute subset of system configurations, or microstates, that stand out in their exceptional stability against the noise introduced by the driving. This principle is clearly supported by observations of SO in 2D granular systems, which evolve toward stable configurations of coordinated cell structures and stresses. Based on this principle, a statistical mechanical formalism has been constructed for SO in those 2D systems and its use outlined. Similar stability measures have been identified in 3D granular systems, which makes possible a straightforward extension to a similar model and its outline has been discussed in detail.
The principle goes beyond granular systems - driven disordered systems always experience noisy dynamics and, when their steady states self-organise, they exclude the overwhelming majority of theoretically possible configurations which do not survive the noise. It also goes beyond the previously proposed principle of minimisation of dissipation energy Cyetal21 ; Keetal23 , which is only a special case that applies to systems where energy is the main driver of the SO. The generality of this principle makes possible an extension of the statistical mechanics approach to describe emergence of SO in wider range of systems. This extension has been formulated here explicitly and it involves identification of quasi-particles; their DOFs; stability measures, which depend on the DOFs; and a survivability function, whose value decreases with the stability measures. The latter is used in a Boltzmann-like exponential, which determines the occurrence probability of system configurations, and a partition function has been constructed. From the partition function, all expectation values of the self-organised state can be derived. This complements the idea in Keetal23 of the system exploring the configuration space in that it parks on configurations that minimise locally the survivability function. The application of the general formalism has been illustrated by constructing a statistical mechanics model for a very simple version of the laning problem.
@@
There are parallels between this formulation and its analogue in thermal systems, in which the microstates are the system-wide configurations and the particles’ energies and potentials are used as stability measures. The thermal survivability function is the free energy, which depends on the quasi-particles’ DOFs and whose lowering increases a microstate’s occurrence probability and its survivability. Following this understanding, some equilibrium macrostates can also be viewed as self-organised. An example is solidification under temperature lowering, which is analogous to increasing $J_{i}$ in (4). The overwhelming number of microstates with very energetic particles are strongly unstable and only the minute crystalline microstates survive to be observed. There are, however, differences. The energy in thermal systems is a constant of the motion, from which dynamic equations for the DOFs can be derived. The survivability function only describes stability in out-of-equilibrium systems and, as far as this author knows, cannot play such a role. Additionally, thermal systems incur indigenous thermal fluctuations while in a-thermal systems it is the driving that gives rise to noise. Nevertheless, the similarities to traditional statistical mechanics are useful as guidelines for constructing general models of SO.
@@
One can argue that the above principle also applies to the SO that is evident in the evolution of biological species. Those systems are driven by many constraints: resources, predators, environmental, etc., which they must adapt to and survive. The dynamics are then extremely noisy because the constraints’ magnitudes and even number can fluctuate wildly. The principle of the survival of the fittest is exactly the same - only elements that can survive these fluctuations, be it cells or larger organisms, remain. While this may mean that, in principle, it could be possible to formulate a statistical mechanical model for species evolution, it is a daunting task due to the enormous number of potential DOFs and stability measures.
@@ Acknowledgement:
Competing interests:
The author declares no competing interests.